### v1

# Keyword-Aware ASR Error Augmentation for Robust Dialogue State Tracking

Abstract
Dialogue State Tracking (DST) is a key part of task-oriented dialogue systems, identifying important information in conversations. However, its accuracy drops significantly in spoken dialogue environments due to named entity errors from Automatic Speech Recognition (ASR) systems. We introduce a simple yet effective data augmentation method that targets those entities to improve the robustness of DST model. Our novel method can control the placement of errors using keyword-highlighted prompts while introducing phonetically similar errors. As a result, our method generated sufficient error patterns on keywords, leading to improved accuracy in noised and low-accuracy ASR environments.
1 Introduction
Figure 1: Illustration of EPA process.
Task-oriented dialogue systems (TODs) assist users in achieving specific objectives through conversations and are used in various sectors, including customer service and hotel reservations. A crucial component of these systems is Dialogue State Tracking (DST), which extracts vital information from conversations in a slot-value format (e.g., hotel-name: Claire Hotel). This information is essential for querying databases and generating responses Young et al. (2013).
However, DST models face significant challenges in spoken dialogue environments, where user utterances are converted into text by automatic speech recognition (ASR). Notably, Soltau et al. (2022) observed a drastic reduction in model accuracy from 41.6% to 23.6% in such environments [^1] Joint goal accuracy result of base size model.. This decline is primarily due to ASR errors, which frequently misrecognize named entities—a key target in DST Nechaev et al. (2021).
To address ASR inaccuracies, data augmentation has emerged as a viable, cost-efficient strategy. Existing text augmentation methods, such as word swapping Wei and Zou (2019) and back translation Sennrich et al. (2015), do not maintain audio similarity with the original text, leading to discrepancies with ASR error patterns. To bridge this gap, Sharma et al. (2020) and Jacqmin et al. (2023) synthesized audio from text with text-to-speech (TTS) model Shen et al. (2018) and processed it through ASR, while Hrinchuk et al. (2020) and Zhang et al. (2021) employed translation model structure to introduce ASR-like errors directly into texts.
Despite these advancements, prior methods often fail to provide sufficient error patterns for DST model training. Accurately identifying key terms is vital for DST performance; thus, models need to be trained on a broad spectrum of ASR-errored keywords. Unfortunately, many current strategies do not ensure that errors are positioned within critical keywords, often generating trivial examples by altering non-essential words such as random words Wei and Zou (2019) or sentence structure Sennrich et al. (2015). This oversight results in sub-optimal DST performance against ASR errors.
To address these limitations, we introduce Error Positioning Augmentation (EPA), a straightforward yet effective method that ensures sufficient errors in keywords. Our method leverages large language models (LLMs) Ouyang et al. (2022); Touvron et al. (2023); Zhang et al. (2022), which have demonstrated impressive capabilities in semantic augmentation Whitehouse et al. (2023); Sahu et al. (2023) and precise text generation control Sun et al. (2023); Liang et al. (2024). Despite their strengths, LLMs’ potential for phonetic augmentation remains largely unexplored.
In our method, we utilize in-context Brown et al. (2020) learning to guide the augmentation process. By providing examples that highlight the key parts, we enable the model to position errors in keywords. Surprisingly, without collecting extensive user voice datasets, a manually created small text dataset (fewer than 10 samples) can generate a wide variety of ASR-errored keywords. This simplifies the error generation process. Additionally, before keyword-focused augmentation, we augmented the text to incorporate utterance-level overall ASR errors, introducing diverse errors from open-source audio datasets. This two-step approach effectively harnesses LLMs for both overall utterance-level and targeted keyword-level phonetic augmentation.
In our experiments, EPA achieved a diversity in targeted keywords (95.4%) and high variance in overall text (1.81x) while maintaining audio similarity (91.5%). During practical DST evaluations, EPA significantly enhanced the robustness of DST models, increasing accuracy from 45.76% to 51.12%. Our analysis indicates that these improvements are primarily due to keyword augmentation, which significantly reduces errors in ASR-errored values.
2 Method
2.1 Notation
Before detailing each step, we first clarify the notation. Dialogue context from turn 1 to $t$ is denoted as $D_{t}$=$\{(s_{1},u_{1}),...,(s_{t},u_{t})\}$ where $s$ denotes for system and $u$ for user utterance. DST model predicts the dialogue state (also called belief state) $B_{t}$ given $D_{t}$. $B_{t}$ is composed with slot $sl$ and value $v$ pairs, denoted as $B_{t}$ = $\{({sl}_{1},v_{1}),...,({sl}_{J},v_{J})\}$ , where ${sl}_{j}$ and $v_{j}$ is $j$-th slot name and value. $J$ is the total number of slots.
Idx Method Examples
1 Original I’d like to find a vegetarian restaurant, if possible.
+EPA I’d hike to find a veggie tarian restroom, if possible.
2 Original I am going to auburn.
+EPA I am flowing to auburng.
3 Original Tuesday, going to bailey’s crossroads please.
+EPA Tuesday, going to baley’s crossroads, peas.
4 Original Hi! Could you find me a train to loris on thursday?
+EPA Oh! Could you find me a trai to lorri on thursdae?
5 Original Ashby is my destination.
+EPA Ashy’s my desity.
Table 1: Examples of ASR errors from EPA. We have highlighted utterance level overall errors in yellow and keyword-specific errors in blue.
2.2 Step 1: ASR Error for Overall Utterance
In this step, we augmented the overall utterance by introducing general ASR errors. We began by constructing example sets for in-context learning, utilizing an open-source audio dataset Ardila et al. (2020). From this dataset, we randomly selected 300 hours of audio along with their corresponding gold transcripts ($g$) and transcribed the audio using an off-the-shelf ASR model (e.g., Whisper-base Radford et al. (2022)) to obtain the erroneous transcriptions ($e$). We denote this example dataset as $DB=\{(g_{1},e_{1}),\ldots,(g_{I},e_{I})\}$.
Next, we augment the overall utterance level of $u$ by guiding the LLM using examples retrieved from $DB$ based on audio similarity (see Figure  1, Step 1). To calculate audio similarity, we converted the characters to phonemes using the International Phonetic Alphabet (IPA) 222 https://pypi.org/project/eng-to-ipa/ and then computed the similarity between the phonemized $u$ and $DB$. After selecting the $k$ examples, we concatenated the instructions, examples, and $u$ into a single prompt and provided it to the LLM. This process resulted in the overall ASR-errored user utterance $\dot{u}$. Concretely, $\dot{u}$ can be obtained by
$\dot{u_{t}}=LLM(Inst_{1}\oplus(g_{1},e_{1})\cdots(g_{k},e_{k})\oplus u_{t}).$ (1)
2.3 Step 2: ASR Error for Keywords
Method Features Low-acc ASR Noised Aud. Paraphrased High-acc ASR
Aud. Utt-aug Key-aug LLM JGA N-Acc JGA N-Acc JGA N-Acc JGA N-Acc
Baseline - - - - 29.88 45.76 29.70 46.77 28.92 48.79 34.87 52.07
AEDA Karimi et al. (2021) - - - 29.90 46.46 29.74 47.48 29.12 48.86 34.94 52.32
EDA Wei and Zou (2019) - - - 29.22 47.65 28.70 49.51 28.08 49.99 33.68 53.78
BT Sennrich et al. (2015) - - - 31.69 49.17 31.26 50.98 29.90 51.73 36.27 54.81
TTS-ASR - - 29.94 46.34 29.99 47.37 29.08 48.88 35.07 52.03
ASR-translation - - 30.40 47.65 30.14 48.45 29.54 50.38 35.25 53.66
[…]
   w/o Keyword Aug - 31.43 49.63 31.51 50.56 30.41 52.02 36.34 54.57
EPA (Llama 7B) 31.54 51.12 31.55 52.27 30.10 53.55 36.22 55.49
   w/o Keyword Aug - 31.12 50.33 31.44 52.07 30.01 53.49 35.70 54.90
EPA (GPT3.5) 32.39 51.12 32.24 52.70 30.95 53.34 36.61 55.87
   w/o Keyword Aug - 31.31 50.67 31.13 52.29 30.06 52.85 35.40 55.80
Olisia Jacqmin et al. (2023) - - - - 30.17 46.25 30.43 48.07 29.13 49.21 36.1 52.58
Table 2: Comparison of various augmentation methods in enhancing the robustness of DST models across different ASR environments. All results were averaged over three seeds for better consistency.
While Step 1 introduces general ASR errors into ${u}$, it does not ensure that these errors enhance the diversity of keywords. To create a more effectivetraining dataset, we generated keyword-focused ASR errors (see Figure 1, Step 2).Here, we manually created several examples from the original DST dataset by explicitly highlighting the keywords in the clean text using the <hl> tag ($\tilde{g}$) and concatenated them with the ASR-errored versions ($\tilde{e}$). We set the $v$ in turn $t$ as keywords. We applied the same highlighting to the inference target $\dot{u}$ and then concatenated the instructions, examples and highlighted $\dot{u}$ into a single prompt, as in Step 1. We perform the in-context learning with this prompt to generate the augmented utterance $\ddot{u}$, containing overall ASR and keyword-specific errors. Explicitly, we get $\ddot{u}$ by
$\ddot{u_{t}}=LLM(Inst_{2}\oplus(\tilde{g}_{0},\tilde{e}_{0})\cdots(\tilde{g}_{k},\tilde{e}_{k})\oplus\dot{u}_{t}).$ (2)
We set the example numbers as 3 in both Step1 and Step2.
After generating the ASR-errored examples, we trained a DST model using both the EPA-generated data and the original seed data. Specifically, we fine-tuned T5-base Roberts et al. (2019) models for DST tasks, with detailed training procedures outlined in the Appendix B.
3 Augmentation Results of EPA
We first assessed the quality of ASR errors simulated by the EPA method to guarantee they closely resembled genuine ASR errors. Table 1 demonstrates that with overall utterance and keyword error augmentation, EPA effectively generates sentences that align with our objectives: maintaining phonetic similarity to the original text while confirming errors in keywords.
4 Experiments
4.1 Experimental Setup
Dataset. The DSTC11 dataset Soltau et al. (2022), an audio version of MultiWOZ 2.1 Eric et al. (2019), comprises 8,000 dialogues for training, 1,000 for validation, and 1,000 for testing. To enhance generalization, we conducted experiments across four distinct ASR environments, characterized by word error rate (WER) [^3] Based on LibriSpeech.test-clean Radford et al. (2022). and noise levels: (1) a low accuracy ASR model (WER > 0.03), (2) a café and traffic noised audio, (3) a paraphrased setting where users naturally paraphrased the transcriptions, and (4) a high accuracy ASR model (WER < 0.03) enviroment.
Metrics. For overall performance evaluation, we used joint goal accuracy (JGA), which requires all slot-value pairs to match the gold label. We also reported named entity accuracy (N-acc), the average accuracy across named entity slots.
LLMs. we primarily used GPT-3.5 Ouyang et al. (2022). Additionally, as this is the first attempt to augment pronunciation-related errors with LLMs, we also included Llama2-7B Touvron et al. (2023) and OPT-6.7B Zhang et al. (2022).
Compared methods. We compared our method with two established approaches: text-based augmentations, AEDA Karimi et al. (2021), EDA Wei and Zou (2019), and Back Translation (BT) Sennrich et al. (2015), and audio-aware augmentation methods, using synthesized audio (TTS-ASR) and translation model structure (ASR-translation). Lastly, we included Olisia Jacqmin et al. (2023), the top-ranked method in the DSTC11 competition (Appendix C).
4.2 Robustness Improvement through EPA
The results in Table 2 demonstrate the effectiveness of EPA in enhancing robustness to ASR errors in the DST task. Remarkably, EPA outperformed existing text-based and audio-based augmentation methods across all environments, showing substantial improvement in N-Acc with gains of over 4%, compared to the modest 1-2% improvements of other methods. This trend is particularly pronounced in challenging environments, such as low-accuracy ASR and noisy audio conditions, and is also observed in LLaMA-7B and OPT 6.7B. The ablation study revealed that keyword error augmentation consistently boosted performance across all LLMs, capitalizing on its utility to optimize DST performance against ASR errors.
4.3 Error Analysis
Method Error Type
Wrong Ignore Spurious
Noised Audio
Baseline
$\triangledown$0%
$\triangledown$0%
$\triangledown$0%
EPA w/o Key-aug
$\triangledown$5.29%
$\triangledown$3.72%
$\triangledown$6.31%
EPA
$\triangledown$8.19%
$\triangledown$7.25%
$\triangledown$1.33%
High-acc ASR (WER<0.03)
Baseline
$\triangledown$0%
$\triangledown$0%
$\triangledown$0%
EPA w/o Key-aug
$\triangledown$1.57%
$\triangledown$3.34%
$\triangledown$6.52%
EPA
$\triangledown$7.16%
$\triangledown$2.43%
$\triangledown$0.75%
Table 3: Ablation study with error analysis. Wrong indicates the model predicts incorrect values, Ignore refers to the model ignored mentioned slots, and Spurious denotes the model predicting values for unmentioned slots. Actual error numbers are in parentheses.
To examine the effects of keyword augmentation in more detail, we conducted ablation studies and analyzed error reduction rates by categorizing errors into three types. Table 3 presents the percentage reduction in error rates compared to the baseline. The results demonstrate that EPA is effective in "Wrong" and "Ignore" error types, and keyword augmentation highly contributed to this improvement by decreasing the error rate from 5.29% to 8.19%. Interestingly, while keyword augmentation led to substantial reductions in the "Wrong" type error, it increased the "Spurious" error type. This suggests that the model may become overly cautious, occasionally overestimating the presence of named entities.
4.4 Qualitative Assessment of EPA Method
Method Uniq. Words NE.chg [ % ] Audio Sim.[%]
Baseline 1 - -
AEDA Karimi et al. (2021) 1.00$\times$ 44.29 91.57
EDA Wei and Zou (2019) 0.86$\times$ 70.03 61.14
BT Sennrich et al. (2015) 1.21$\times$ 73.46 77.17
TTS-ASR 1.01$\times$ 38.84 98.93
Translating 0.84$\times$ 39.59 94.07
EPA 1.81$\times$ 95.47 91.57
   w/o Keyword Err. 1.57$\times$ 68.81 93.14
Table 4: Assessment of EPA dataset quality across three key aspects (Unique word increased rate(Uniq. Words, Named entity changed rate (NE.chg), and audio similarity (Audio Sim.).
To evaluate the word diversity and audio similarity of the augmented text, we performed a qualitative assessment in Table 4. This assessment includes metrics such as the unique word increase rate, named entity change rate, and audio similarity  [^4] Normalized phoneme edit distance is used for audio similarity calculating.. The results demonstrate that EPA achieves remarkable diversity in unique words (1.81x) and the highest rate of named entity augmentation (95.47%), while maintaining considerable audio similarity (91.57%). Furthermore, keyword augmentation significantly contributes to the improvement, particularly in the named entity change rate (from 68.81% to 91.57%), ensuring that the errors exhibit sufficiently diverse patterns in keywords.
5 Conclusion
We propose a novel data augmentation method tailored for DST tasks that ensures sufficient error patterns in both key phrases and overall text. By leveraging LLMs for their controlled text generation capabilities, we strategically place errors within key phrases. Our method demonstrates substantially improved robustness in DST by generating diverse, plausible keyword errors. Error case analysis reveals that keyword augmentation significantly enhances robustness against ASR errors. As the pioneering research in leveraging LLMs for generating ASR errors, we hope this work lays a strong foundation for future phonetic-based augmentation research.
Limitations
First, the introduction of keyword augmentation, while reducing "Wrong" and "Ignore" errors, resulted in an increase in "Spurious" errors, indicating a potential trade-off in error types. Additionally, our experiments were conducted using only the T5-base model. However, we believe that our method will work effectively with other models as well. Furthermore, we did not explore the application of our method to other similar datasets, such as those used in spoken language understanding (SLU) or named entity recognition (NER) tasks. Future work should consider these areas to validate and extend the applicability of our approach.
Appendix A Baseline Performance Comparison with Clean Text
For comparison, we report the baseline performance on an error-free, clean test dataset. Please note that DSTC11 Soltau et al. (2022) does not provide a text script for the test dataset, so we are manually cleaning 50 dialogues to ensure they are error-free. In the experiment, the baseline model achieved a JGA score of 45.2 % and an N-ACC score of 86.5 % in an ASR error-free environment. Compared to the JGA, which is 34.8 %, and N-ACC, which is 52.07 %, in the ASR-errored environment (High-acc ASR model environment), this discrepancy highlights the significant impact of ASR errors on performance degradation.
Appendix B Training Details
In training models, we used T5-base Roberts et al. (2019) as the backbone model and instructed the model to generate the $B_{t}$ by given $D_{t}$ in sequence to sequence manner, as in Su et al. (2021) and the loss function is
$L=-\sum_{t=1}^{T}\log P(B_{t}|\text{Inst},D_{t}).$ (3)
We set the learning rate as 4e-5 and used the AdamW Loshchilov and Hutter (2017) optimizer. One GeForce RTX 3090 is used for training and the batch size is 16. Trained until reaching the max patient, which is 3.
Appendix C Experiment Details
C.1 Dataset
  • Low-acc ASR environment: Whisper-base model (74M)Radford et al. (2022) is used for transcription. WER on LibriSpeech.test-clean is 0.05.
  • Noisy audio environment: Incorporated authentic cafe and traffic noise from https://freesound.org/ with a 10 to 20 Signal-to-Noise Ratio (SNR) and transcribed it using the Whisper large model.
  • Paraphrased environment: When recording the audio, the text was paraphrased to more closely resemble spoken language Soltau et al. (2022).
  • High-acc ASR environment: Whisper-large model (1550M) is used for transcription. WER on LibriSpeech.test-clean is 0.027.
C.2 Comparison Methods
  • AEDA Karimi et al. (2021): We randomly inserted punctuation marks, effectively maintaining the original word order.
  • EDA Wei and Zou (2019): We augmented data by applying edit-based technique that implements four rule-based modifications—synonym replacement, random insertion, swapping, and deletion.
  • Back Translation Sennrich et al. (2015): We translated original texts to error texts and then back to the original texts for generating syntactic variations during the process. We use English to German 555facebook/wmt19-en-de and German to English666facebook/wmt19-de-en models as translator.
  • TTS-ASR : We used Tacotron2 Shen et al. (2018) for the TTS model to synthesize the audio and use Whisper-base Radford et al. (2022) as an ASR model to simulate the ASR errors.
  • ASR translation: We employed a sequence-to-sequence structure to translate clean text into ASR-errored text. Our training set comprised 300 hours of paired clean and ASR-errored text, as detailed in Section 2.2. We fine-tuned the model based on the T5-base architecture Roberts et al. (2019), using the loss function defined in equation 4. The loss function is as follows:
    $L=-\sum_{i=1}^{I}\log P(e_{i}|g_{i}).$ (4)
Appendix D Prompt Example
Step 1 Prompt Example
Generate ASR error augmented text with similar pronunciation but different words based on the given gold text examples.
Apply character and word substitutions, additions, or deletions while maintaining the overall pronunciation and context.
Error rate should be high
{Example 1}
Original: they have a single naupliar eye
ASR-errored: they have a single nor pure eye
{Example 2}
Original: i must have saint louis then huzza
ASR-errored: i must have st louis then hazard
{Example 3}
Original: i wonder uncle did not have her come
ASR-errored: i wonder uncle did not have a problem
{Example 4}
Original: i don’t want to have to pay for parking.
ASR-errored:
Step 2 Prompt Example
Change the key words in <hl> tag, to having a ASR error. ASR error has similar pronounciation with the correct word, but different charater.
Here is some example
{Example 1}
Original: I want to buy a book about <hl>luwombo best</hl> restaurant. Keywords : luwombo best
Result: I want to buy a book about luwambo vest restaurant.
{Example 2}
Original: hi, i’m looking for a bus that is depart from <hl>eliot<hl/> and arriving to <hl>holiday inn williamsport<hl/>? Keywords : eliot, holiday inn williamsport
Result: hi, i’m looking for a bus that is depart from Ellyot and arriving to holliday inn william’s port
{Example 3}
Original: the <hl>chabuton ramen<hl/> is a restaurant on the east. Keywords : chabuton ramen
Result: the shabuton raymond is a restaurant on the east.
{Example 4}
Original: i was hoping to see local faces while in cambridge. some <hl>entertainment</hl> would grate. Keywords: entertainment
Result:
Method Low-acc ASR Noised Aud. Paraphrased High-acc ASR
JGA N-acc JGA N-acc JGA N-acc JGA N-acc
Baseline 30.05 46.48 29.80 47.49 29.08 48.85 34.73 52.31
AEDA 29.99 46.53 29.80 47.60 28.95 49.02 34.94 52.40
EDA 29.16 47.67 28.93 49.50 28.12 50.10 33.74 53.66
BT 31.54 49.21 31.43 51.25 29.90 51.60 36.25 54.85
TTS-ASR 30.09 46.16 30.32 47.74 29.26 49.31 35.28 51.95
FT 29.98 47.70 29.95 48.49 29.72 50.43 34.82 53.35
EPA (GPT3.5) 32.56 51.62 32.27 53.48 31.10 53.91 36.65 56.14
Baseline 29.82 45.63 29.75 46.20 28.55 48.75 34.83 51.44
AEDA 29.77 46.29 29.69 47.40 29.26 48.93 34.93 52.26
EDA 29.27 47.48 28.57 49.41 28.20 49.99 33.59 53.68
BT 31.66 49.16 31.05 50.81 29.88 51.75 36.19 54.73
TTS-ASR 29.75 45.99 29.80 46.87 29.10 48.12 34.97 51.49
FT 30.86 47.47 30.52 48.53 29.63 50.11 35.73 53.59
EPA (GPT3.5) 32.33 50.74 32.41 52.28 31.02 53.15 36.75 55.82
Baseline 29.77 45.18 29.56 46.61 29.12 48.76 35.05 52.47
AEDA 29.94 46.55 29.72 47.45 29.16 48.62 34.94 52.29
EDA 29.22 47.79 28.59 49.61 27.92 49.88 33.71 53.99
BT 31.86 49.13 31.31 50.87 29.91 51.85 36.38 54.85
TTS-ASR 29.98 46.86 29.84 47.51 28.88 49.20 34.97 52.66
FT 30.37 47.78 29.94 48.32 29.26 50.60 35.20 54.05
EPA (GPT3.5) 32.27 51.01 32.04 52.33 30.72 52.96 36.42 55.66
Table 5: Experiment result with different seeds.